OpenStudy (anonymous):

if 10x198=(10xm)-(10x2) what is the value of m?

OpenStudy (asnaseer):

can you see a common factor on the right-hand-side of the equals sign?

OpenStudy (anonymous):

no new to this

OpenStudy (asnaseer):

what number appears in both terms on the right-hand-side?

OpenStudy (anonymous):

this is the way question is

OpenStudy (asnaseer):

I know, I am asking you a question in order to help guide you to the answer. so, can you spot what number appears in both terms on the right-hand-side?

OpenStudy (anonymous):

10 right

OpenStudy (asnaseer):

good

OpenStudy (asnaseer):

so now we can factor out the 10 and re-write the equation as follows:\[10\times198=10\times m-10\times2=10\times(m-2)\]

OpenStudy (asnaseer):

do you understand so far?

OpenStudy (anonymous):

little confusing

OpenStudy (asnaseer):

it might be easier to use an example to illustrate this... \[10\times3-10\times2=30-20=10\]but this could also have been done in this way:\[10\times3-10\times2=10\times(3-2)=10\times1=10\]

OpenStudy (asnaseer):

so, in general:\[10\times a-10\times b=10\times(a-b)\]

OpenStudy (asnaseer):

ok, so are you ready for the next bit?

OpenStudy (asnaseer):

we got to this point:\[10\times198=10\times m-10\times2=10\times(m-2)\]so we now know:\[10\times198=10\times(m-2)\]next, notice that both left-hand-side and right-hand-side have something multiplied by 10. so we can divide both sides by 10

OpenStudy (anonymous):

my brain is going to pop thanks for your help.

OpenStudy (asnaseer):

np - just one more step to go and you'll have the answer.

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

so, if we divide both sides by 10, we get left with:\[198=m-2\]

OpenStudy (asnaseer):

now, can you see what value m must be?

OpenStudy (asnaseer):

"something" minus 2 is 198, so that "something" must be equal to?

OpenStudy (anonymous):

196

OpenStudy (asnaseer):

is 196 - 2 = 198?

OpenStudy (anonymous):

200

OpenStudy (asnaseer):

perfect! you got it! :)

OpenStudy (asnaseer):

I would suggest you practice with lots of problems like this, I'm sure you'll soon get the hang of them

OpenStudy (anonymous):

ok got to go are we done?

OpenStudy (asnaseer):

yes :)

OpenStudy (anonymous):

cool thanxs

OpenStudy (asnaseer):

yw :)